# Yeah, Un, Oh: Continuous and Real-time Backchannel Prediction with Fine-tuning of Voice Activity Projection

### v1

@@ 1 Introduction
Note that the source codes and trained models are publicly available  [^1] https://github.com/inokoj/VAP-Realtime.
@@ 3.1 Attentive Listening Dataset
We have collected spoken dialogue data using a Wizard-of-Oz (WOZ) setup. In this experiment, the android ERICA Inoue et al. (2016) was employed, with a human operator remotely controlling it, which was transmitted and played through ERICA’s speaker system (Figure 2). The dialogue task focused on attentive listening, where human participants shared personal experiences, and ERICA actively engaged as a listener Inoue et al. (2020b); Lala et al. (2017). This task was advantageous because it allowed for the collection of numerous backchannel responses by ERICA. The participants comprised two demographic groups: students and older adults. Each group was provided with a prompt to guide their conversation; for instance, students discussed “challenges during the COVID-19 pandemic,” while the elderly participants reflected on “memorable travel experiences and recent favorite meals.” ERICA’s operators were three experienced actresses, with one participating in each session. In total, we recorded 109 dialogue sessions, each lasting approximately 7 to 8 minutes. The data were randomly divided into 87, 11, and 11 sessions for training, validation, and testing purposes, respectively. We subsequently transcribed the dialogues and annotated ERICA’s backchannel responses.
@@ 5.1 Voice Activity Projection
The VAP model employed in this study is constructed upon a Transformer-based architecture designed to emulate human-like predictive capabilities. As illustrated in Figure 4, the architecture of the VAP model processes stereo audio signals originating from two participants [^2] A detailed explanation of the model’s architecture is provided in a previous work Inoue et al. (2024a). Note that this model integrates the listener’s audio as one of the input channels. This allows self-generated backchannels to be fed back into the model, distinguishing it from other existing models. This functionality plays a crucial role in preventing multiple consecutive backchannels, which may appear unnatural.
@@
The concatenated vector is subsequently passed through linear layers for two distinct tasks: voice activity projection and voice activity detection. The primary task, voice activity projection, is represented as a 256-dimensional vector that predicts the future two-second voice activity of the two participants Ekstedt and Skantze (2022b). In contrast, the voice activity detection task focuses on the current frame, generating two binary vectors that represent the present voice activity.
@@ 6.1 Timing Prediction
The first experiment focuses on the initial tasks outlined in Section 4.1, which involve a binary classification of backchannels or non-backchannels. We prepared several comparative methods, including random classification (always outputs positive), as detailed below: (i) Baseline consists solely of the audio encoder CPC and a linear head. While other methods in this study freeze the CPC during training, this baseline approach fine-tunes the CPC model itself. As the two input audio channels are separately fed into the CPC, their distinct output vectors are concatenated and then passed to the linear head. (ii) ST w/o PT refers to the single-task (ST) model, where the loss function only includes $L_{bc}$ from Equation (1). Moreover, this model does not involve any pre-training (PT) of the VAP model. (iii) ST w/ PT introduces pre-training of the VAP model in addition to the single-task learning. (iv) MT w/ PT represents the proposed method, which incorporates both multi-task (MT) learning, as described in Equation 1, and the pre-training of the VAP model. The evaluation metrics used are F1-score, precision, and recall, calculated in a frame-wise manner, and the F1-score is the most priority indicator. For the hyperparameters in Equation (1), to emphasize the importance of backchannel prediction, we empirically set them as $\alpha=1$, $\beta=1$, and $\gamma=5$. Additionally, to mitigate the impact of the imbalanced dataset, we adjusted the loss weight, assigning a weight five times larger to positive (backchannel) samples compared to negative (non-backchannel) samples.
Table 1 presents a summary of the results for this task. Firstly, all trained methods demonstrated significantly highter scores when compared to the random approach. In particular, the proposed method achieved the highest scores in both the F1 score and precision metrics. These findings suggest that both pre-training and multi-task learning play a crucial role in improving backchannel prediction performance, indicating that this task requires a more robust approach than conventional target-specific training or fine-tuning. Moreover, the VAP model, along with its original loss function, exhibits better generalizability and applicability to other non-linguistic behavior predictions, such as the current backchannel prediction task.
@@
6.2 Timing and Type Prediction
The second experiment involves the prediction of backchannel types, as outlined in Section 4.2. We employed the same comparative methods as in the previous experiment, but adjusted the output dimension of the linear head from 2 to 3 to accommodate the classification of continuers, assessments, and non-backchannels. The evaluation metric remained unchanged; however, we conducted separate evaluations for continuers and assessments.
Table 2 and Table 3 present the outcomes of this task. As with the previous results, both tables demonstrate that the combination of multi-task learning and pre-training significantly enhanced performance, with the proposed method achieving the highest F1-score. When comparing these two types, as anticipated, the prediction of assessment backchannels yielded lower scores. While random prediction offers no meaningful insight, the proposed method exceeded an F1-score of 30. Compared to a recent similar study Choi et al. (2024), the model demonstrated a higher and more promising score. However, given the variations in the datasets employed and the differences in model sizes, a direct comparison under identical conditions is challenging. Nevertheless, such comparisons should be pursued in future research.

### v2

@@ Abstract
 [^0] This paper has been accepted for presentation at the main conference of 2025 Annual Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics (NAACL 2025) and represents the author’s version of the work.
@@ 1 Introduction
Note that the source codes and trained models are publicly available  [^1] We will put the link here after the reviewing process..
@@ 3.1 Attentive Listening Dataset
We have collected spoken dialogue data using a Wizard-of-Oz (WOZ) setup. In this experiment, the android ERICA Inoue et al. (2016) was employed, with a human operator remotely controlling it, which was transmitted and played through ERICA’s speaker system (Figure 2).
The dialogue task focused on attentive listening, where human participants (speaker) shared personal experiences, and ERICA actively engaged as a listener Inoue et al. (2020b); Lala et al. (2017). This task was advantageous because it allowed for the collection of numerous backchannel responses by ERICA. The participants comprised two demographic groups: students and older adults. Each group was provided with a prompt to guide their conversation; for instance, students discussed “challenges during the COVID-19 pandemic,” while the elderly participants reflected on “memorable travel experiences and recent favorite meals.”
ERICA’s operators were three actresses, who had experiences of our past attentive listening dialogue experiments. While backchannel behaviors can be subjective and vary among individuals, the few operators were selected for their experience, ensuring the collection of high-quality backchannel data. Furthermore, the operators participated in a sufficient number of dialogue sessions in this experiment, ensuring both the quantity and quality of the training data, despite the varying speakers in each session.
In total, we recorded 109 dialogue sessions, each lasting approximately 7 to 8 minutes. The data were randomly divided into 87, 11, and 11 sessions for training, validation, and testing purposes, respectively. We subsequently transcribed the dialogues and annotated ERICA’s backchannel responses.
@@ 5.1 Voice Activity Projection
The VAP model employed in this study is constructed upon a Transformer-based architecture designed to emulate human-like predictive capabilities. As illustrated in Figure 4, the architecture of the VAP model processes stereo audio signals originating from two participants [^2] A detailed explanation of the model’s architecture is provided in a previous work Inoue et al. (2024a), consistent with the operation of full-duplex spoken dialogue systems. Note that this model integrates the listener’s audio as one of the input channels. This allows self-generated backchannels to be fed back into the model, distinguishing it from other existing models. This functionality plays a crucial role in preventing multiple consecutive backchannels, which may appear unnatural.
@@
Figure 5: Definition of the VAP state
The concatenated vector is subsequently processed by linear layers for two distinct tasks: voice activity projection (VAP) and voice activity detection (VAD). The primary task, voice activity projection, yields a 256-dimensional state vector that predicts the voice activity of the two participants over the next two seconds Ekstedt and Skantze (2022b), as illustrated in Figure 5. This two-second period is divided into four time intervals: 0-200 ms, 200-600 ms, 600-1200 ms, and 1200-2000 ms. Consequently, there are eight binary bins in total, four for each participant, resulting in 256 ($=2^{8}$) possible combinations of speaking/non-speaking states for each participant within these bins. The voice activity detection task, conversely, focuses on the current frame, producing two binary vectors representing the present voice activity of each participant. Both tasks are trained using the cross-entropy loss function, denoted as $L_{vap}$ and $L_{vad}$ in the following subsection.
@@ 6.1 Timing Prediction
The first experiment focuses on the initial tasks outlined in Section 4.1, which involve a binary classification of backchannels or non-backchannels. We prepared several comparative methods, including random classification (always outputs positive), as detailed below:
  • (i)
    Baseline consists solely of the audio encoder CPC and a linear head. While other methods in this study freeze the CPC during training, this baseline approach fine-tunes the CPC model itself. As the two input audio channels are separately fed into the CPC, their distinct output vectors are concatenated and then passed to the linear head.
  • (ii)
    Zero-shot is based on the direct use of the VAP output proposed in the original VAP work Ekstedt and Skantze (2022b). It is a normalized value obtained by adding the probability of immediate voice activity of the listener (system) (0.0 to 1.2 seconds) and the probability of slightly later voice activity of the speaker (user) (1.2 to 2.0 seconds). A higher value indicates that the immediate system utterance would be short, corresponding to a backchannel.
  • (iii)
    ST w/o PT refers to the single-task (ST) model, where the loss function only includes $L_{bc}$ from Equation (1). Moreover, this model does not involve any pre-training (PT) of the VAP model.
  • (iv)
    ST w/ PT introduces pre-training of the VAP model in addition to the single-task learning.
  • (v)
    MT-ASR performs another type of multi-task learning, where backchannel prediction is trained together with automatic speech recognition, inspired by a recent backchannel prediction model Choi et al. (2024). This subtask was trained to recognize phonemes (19 phones) using the CTC loss function.
  • (vi)
    MT w/ PT represents the proposed method, which incorporates both multi-task (MT) learning, as described in Equation 1, and the pre-training of the VAP model.
The evaluation metrics used are F1-score, precision, and recall, calculated in a frame-wise manner, and the F1-score is the most priority indicator. For the hyperparameters in Equation (1), to emphasize the importance of backchannel prediction, we empirically set them as $\alpha=1$, $\beta=1$, and $\gamma=5$. Additionally, to mitigate the impact of the imbalanced dataset, we adjusted the loss weight, assigning a weight five times larger to positive (backchannel) samples compared to negative (non-backchannel) samples.
@@
Table 1 presents a summary of the results for this task. Firstly, all trained methods demonstrated significantly higher scores when compared to the random and zero-shot approaches. In particular, the proposed method (MT w/ PT) achieved the highest scores in both the F1 score and precision metrics. These findings suggest that both pre-training and multi-task learning play a crucial role in improving backchannel prediction performance, indicating that this task requires a more robust approach than conventional target-specific training or fine-tuning. Moreover, the VAP model, along with its original loss function, exhibits better generalizability and applicability to other non-linguistic behavior predictions, such as the current backchannel prediction task.
@@
6.2 Timing and Type Prediction
The second experiment involves the prediction of backchannel types, as outlined in Section 4.2. We employed the same comparative methods as in the previous experiment, but adjusted the output dimension of the linear head from 2 to 3 to accommodate the classification of continuers, assessments, and non-backchannels. The evaluation metric remained unchanged; however, we conducted separate evaluations for continuers and assessments. Note that the zero-shot approach (defined in Section 6.1) was not applicable to this task and was therefore excluded.
Table 2 and Table 3 present the outcomes of this task. As with the previous results, both tables demonstrate that the combination of multi-task learning and pre-training significantly enhanced performance, with the proposed method (MT w/ PT) achieving the highest F1-score. When comparing these two types, as anticipated, the prediction of assessment backchannels yielded lower scores. While random prediction offers no meaningful insight, the proposed method exceeded an F1-score of 30.